# Improving U-Net Confidence on TEM Image Data with L2-Regularization, Transfer Learning, and Deep Fine-Tuning

### v1

@@ Abstract
With ever-increasing data volumes, it is essential to develop automated approaches for identifying nanoscale defects in transmission electron microscopy (TEM) images. However, compared to features in conventional photographs, nanoscale defects in TEM images exhibit far greater variation due to the complex contrast mechanisms and intricate defect structures. These challenges often result in much less labeled data and higher rates of annotation errors, posing significant obstacles to improving machine learning model performance for TEM image analysis. To address these limitations, we examined transfer learning by leveraging large, pre-trained models used for natural images.
We demonstrated that by using the pre-trained encoder and L2-regularization, semantically complex features are ignored in favor of simpler, more reliable cues, substantially improving the model performance. However, this improvement cannot be captured by conventional evaluation metrics such as F1-score, which can be skewed by human annotation errors treated as “ground truth”. Instead, we introduced novel evaluation metrics that are independent of the annotation accuracy. Using grain boundary detection in UO2 TEM images as a case study, we found that our approach led to a 57% improvement in defect detection rate, which is a robust and holistic measure of model performance on the TEM dataset used in this work. Finally, we showed that model self-confidence is only achieved through transfer learning and fine-tuning of very deep layers.
@@ 1 Introduction
Nanoscale defects, such as grain boundaries, precipitates, and dislocations, play a critical role in controlling the properties and functionality of solid-state materials. Transmission electron microscopy (TEM) has become an irreplaceable tool for investigating these defects, owing to its ultrahigh spatial resolution (sub-Angstrom) [16]. Furthermore, recent advances in faster electron detection and data processing have enabled a big-data approach to characterization techniques such as in-situ TEM and 4D-scanning transmission electron microscopy (4D-STEM) [8]. Terabytes of data can be created in a single hour during 4D-STEM or in-situ TEM experiments [5]. However, extracting meaningful insights from the data has quickly become an enormous bottleneck, since the traditional method of manual image analysis is time-consuming, subject to human bias, and cannot scale with the growing data volume [11]. Therefore, developing high-quality automated approaches for TEM image analysis is of paramount importance.
Since the discovery of convolutional neural networks (CNN), machine learning (ML) models have been able to outperform not only traditional computer vision techniques, but even human abilities on certain image analysis tasks [7]. In the context of TEM, binary segmentation utilizing models from the U-Net family enables pixel-level classification of defect structures, which is especially important for identifying continuous defects like grain boundaries and phase interfaces [10]. However, comparing applications to more typical datasets, the performance of CNN models for TEM image analysis remains inferior. For instance, most of the literature involving U-Net based models, including the seminal paper [9] and popular derivatives like U-Net++ [21], is focused on medical imaging [9]. With these datasets, it is very common to see F1-scores in the range 0.85–0.95. Similar performance can also be seen in applications like forestry [14], crack detection [6], satellite imaging [2], and plant disease [15]. TEM applications, however, usually report lower F1-scores in range 0.5–0.8 [1].
@@
To address these challenges, we explored the use of transfer learning by leveraging advanced ML models pre-trained on large datasets for use with TEM analysis. In particular, we investigated promising combinations of a pre-trained EfficientNet encoder with a U-Net++ decoder, which helped achieve the best performance in 2023 and 2024 for edge detection in regular optical image datasets [19][20]. We also introduced two novel metrics, prediction certainty and prediction abundance, describing the ability of a model to make predictions with high class probabilities. Together, they define a model’s self-confidence and are independent of ground truth accuracy. We demonstrated that U-Net performance can be substantially improved using a pre-trained encoder, fine-tuning of deep layers, and L2-regularization. Although F1-scores remained limited due to ground truth flaws and uncertainty, improved model self-confidence led to a 57% increase in the number of detected defects for the dataset tested, significantly enhancing the accuracy of defect statistics.
As a practical example, we applied our workflow to a TEM image dataset of nanocrystalline UO2 samples used to study grain growth as function of temperature and heavy ion irradiation dose. UO2 is the primary fuel used in current nuclear reactors, and many of its key material properties, such as thermal conductivity, fission gas retention, and fracture toughness, are governed by grain size. Therefore, it is critical to establish reliable correlations between grain size of UO2 and its irradiation conditions for predictive modeling of nuclear fuel performance. The accuracy of said correlations principally depends on the quality of grain statistics, which in turn requires large and representative datasets. As such, hundreds of TEM images like those in Fig. 1 were collected at various dose levels and temperatures. Processing these images by hand is infeasible, so U-Net models were used to automate segmentation of grain boundaries.
@@ 2.1 Grain Boundary Dataset
The dataset used in this work consists of bright-field (BF) TEM images of nanocrystalline thin-film UO2 samples. Details about fabrication, in-situ irradiation, and TEM imaging can be found in [18] [17]. The training dataset was created by selecting 14 images from a variety of experimental conditions and resizing them from 4096$\times$4096 to 1024$\times$1024. The grain boundaries were then labeled using Inkscape with a thickness of 5 pixels. Pixels belonging to a grain boundary were given a value of 1, while pixels belonging to the grain interior were given a value of 0. A one-to-one mapping is ultimately constructed between TEM and annotated images. Examples can be seen in Fig. 1.
Small dataset size can lead to significant variation in model performance when different validation and test hold-out sets are used. To improve their representation of the training set, each image was quartered into four 512$\times$512 images, and 5-fold cross-validation was used. This divides the dataset into five non-overlapping validation sets of equal size, training a model for each one with the same hyperparameters. Statistics can then be calculated for performance metrics, providing a more robust estimate. This also means a test set is not necessary, further improving validation set representation. To improve model performance, training images were augmented eightfold using 90° rotations and flipping over horizontal and vertical axes. Each model thus used a different combination of 360 training images and 11 validation images.
@@ 2.2 Model Architectures and Hyperparameters
To improve U-Net segmentation performance, the encoder is replaced with EfficientNetB7, the largest model from the EfficientNet-v1 family [13], and the decoder has been replaced with U-Net++. Most promisingly, the architecture EfficientNetB7-U-Net++ has been used as the base network of state-of-the-art edge detectors UAED [19] and MUGE [20], which are applicable to grain boundary segmentation. Furthermore, EfficientNetB7 has been pre-trained on ImageNet [4] to ease training and boost encoder performance for small datasets. This is the sole focus of Sec. 4, where the benefits of transfer learning and fine-tuning are evaluated. EfficientNetB7 and U-Net++ were also isolated to show their respective contributions, leading to the four architectures investigated in this study, as summarized in Fig. 3: UNet (b), UNet++ (c), Eff-UNet (d), and Eff-UNet++ (e).
The Adam optimizer was used with a global learning rate of 1e-4 and a batch size of 4. For the yellow convolution blocks in Fig. 2b-e, each Conv2D layer learned $2^{(i+6)}$ filters, where $i=0,1,2,3,4$ refers to the encoder resolution level. ReLU activation was used for all layers except the output layer which used Sigmoid activation. To suppress overfitting, L2-regularization was used to penalize large weights in favor of a more distributed network. The total loss can be written symbolically as
$\mathscr{L}=\textrm{BCE}(\textrm{prediction},\textrm{annotation})+\lambda||\mathbf{w}||^{2},$
where BCE is binary cross-entropy loss, commonly used in binary segmentation, $\mathbf{w}$ is a vector of model weights, and $\lambda$ is the L2-regularization strength. In preliminary testing, it was found $\lambda$ can have substantial effects on model output, specifically on model self-confidence. Given that the architectures in Fig. 2 have different sizes and connections, and thus very different $\mathbf{w}$, a coarse grid search was performed to determine the optimal $\lambda$. This grid search provided the basis for model comparison.
@@ 2.3 Performance Metrics
To evaluate model performance, classification metrics precision, recall, and F1-score were used with a binarization threshold of 0.5. This minimizes precision-recall tradeoff and maximizes F1-score. Two novel metrics, prediction certainty and abundance, were also used to evaluate model self-confidence. As shown in Fig. 3, they operate on raw prediction image histograms, where certainty measures the ability of a model to predict high class probabilities, and abundance measures the proportion of said pixels that are positive predictions. A model with high certainty tends to avoid uncertain predictions where a pixel could belong to either a grain boundary or non-grain boundary with nearly equal probability. A model with high abundance tends to predict more grain boundaries overall, which is important for an imbalanced dataset. We describe model self-confidence comparatively as the possession of both certainty and abundance, which, importantly, do not rely on the existence of a ground truth.
The final metric, grain detection rate, was used to provide a holistic measure of model usefulness. It implicitly depends on all other metrics evaluated, as a model must have good classification performance and self-confidence to predict well-defined grains that can be detected by a post-processing algorithm. Models with the highest grain detection rate provide the most accurate grain statistics, and are therefore of supreme practical value. To detect grains from segmented TEM images, we used an algorithm developed in previous work called Convex Hull Approximate Contour (CHAC). CHAC is based on thermodynamics principles stating that grain boundaries should be convex with respect to each other, so only closed convex contours in the processed TEM images should be counted as grains.
@@ 3.1 Visual Comparison of Predictions
Initial insights into model performance can be gained by examining raw model predictions visually. A representative validation prediction is shown for all architectures and L2-regularization strengths in Fig. 4. When no regularization was used, all architectures appeared to perform similarly. However, upon increasing $\lambda$, UNet and UNet++ responded quite differently than Eff-UNet and Eff-UNet++. The latter architectures exhibited greater tolerance to regularization, as over-regularization occurred at much higher $\lambda$ values. Furthermore, predictions made near $\lambda$=1e-3 demonstrated significant quality improvements, where models clearly predict higher class probabilities, and thus have greater certainty. This has the primary benefit of reducing grain boundary noise in binarized predictions, as pixel values tend to be far from the threshold (0.5), and are thus not left up to meaningless noise. Finally, unlike changing the encoder, there was no obvious difference between using U-Net and U-Net++ decoders.
3.2 Model Architectures and Hyperparameters
@@
Fig. 5 summarizes the six metrics used to evaluate model architectures for different values of $\lambda$. Like the visual results in Fig. 4, architectures behave similarly when $\lambda$=0, other than Eff-UNet and Eff-UNet++ having marginally lower precision and recall. As $\lambda$ is increased, recall, certainty, abundance, and grain detection rate reached their maximum at $\lambda$=1e-4 for UNet and UNet++, and between $\lambda$=5e-4 and $\lambda$=1e-3 for Eff-UNet and Eff-UNet++. Coincidentally, precision reached its minimum at these points. Furthermore, the following observations can be made: (1) F1-score showed little-to-no improvement despite considerable improvement to grain detection rate, (2) there is a large divergence between U-Net and EfficientNet encoder architectures, with the latter showing much greater certainty, abundance, and grain detection rates in Figs. 5d-f, and (3) given the width of the error bars, there is no practical benefit from using the U-Net++ decoder.
@@
Most significantly, by using EfficientNet and L2-regularization, grain detection rate increased by 57% over unregularized UNet. For the 14 TEM images in this dataset, that is a difference of 2000 grains, or about 150 more grains detected per image. This is clearly captured by certainty and abundance, but the classification metrics in Figs. 5a-c do not explicitly show this, other than a 7% improvement to recall. This highlights the value of model self-confidence and holistic metrics like grain detection rate. As discussed next in Sec. 3.3, the tradeoff between precision and recall as $\lambda$ increases does provide some insight, but F1-score may not be entirely reliable for images that have annotation errors.
@@
Figure 5: Metrics obtained from 5-fold cross-validation for $\lambda$ optimization. Figs. a-c are classification metrics precision, recall, and F1-score respectively. Figs. d-e are proposed metrics certainty, abundance, and grain detection rate respectively. Error bars correspond to ±$\pm 1\sigma$
@@ 3.3 Importance of L2-Regularization
As shown in Fig. 5c, the F1-score is limited to 0.75, regardless of architecture or regularization strength. To investigate this limitation, classification performance was visualized in Fig. 6 using a technique we call Confusion Matrix Color Mapping (CM2). Here, every pixel in a binarized prediction image is colored according to its classification outcome in the binary confusion matrix, i.e., green for true positive, blue for false positive, and red for false negative pixels. Importantly, orange arrows indicate examples of grain boundaries which were well-supported by the TEM image and were detected by the Eff-UNet model, but not present in the annotation. In other words, the human annotator missed these grain boundaries. Annotation errors act as a significant limitation of classification performance where good model predictions are incorrectly penalized, or where any attempt is bound to disagree.
To minimize loss, a model attempts to learn a set of operations that replicate the annotated grain boundaries like those in the middle column of Fig. 6. However, when true grain boundaries are mistakenly missing in the annotation, the model cannot rationalize these inconsistencies, potentially increasing model self-uncertainty. Any attempt to exactly fit these errors acts as a contradictory force to its typically well-supported operations. If annotation errors are interpreted by the model as high-level features with complex semantic content, the model could begin to overfit by trying to learn these features. The benefits of L2-regularization then become obvious: by constraining model complexity, it discourages fitting such annotation noise, and can instead help the model prioritize simpler, more reliable cues, thus reducing model self-uncertainty and promoting model self-confidence.
The theory above is supported well quantitatively. As shown in Fig. 5, without regularization all four architectures have relatively high precision and low recall, meaning they guess less but are usually correct, and have very low prediction certainty and abundance. This indicates that they are skeptical. However, applying optimal L2-regularization, the architectures shift toward low precision and high recall, where they now guess more often and are less concerned with correctness. Certainty and abundance also increase substantially, suggesting less influence from human errors and more confidence with their predictions. Among all four architectures, Eff-UNet and Eff-UNet++ show the most significant improvement to prediction quality with grain detection rate reaching the highest observed levels, likely because of the greater capacity and redundancy supplied by EfficientNetB7.

### v2

@@ Abstract
With ever-increasing data volumes, it is essential to develop automated approaches for identifying nanoscale defects in transmission electron microscopy (TEM) images. However, compared to features in conventional photographs, nanoscale defects in TEM images exhibit far greater variation due to the complex contrast mechanisms and intricate defect structures. These challenges often result in much less labelled data and higher rates of annotation errors, posing significant obstacles to improving machine learning model performance for TEM image analysis. To address these limitations, we examined transferring learning by leveraging large, pre-trained models used for natural images.
We demonstrated that by using the pre-trained encoder and L2-regularization, semantically complex features are ignored in favor of simpler, more reliable cues, substantially improving the model performance. However, this improvement cannot be captured by conventional evaluation metrics such as F1-score, which can be skewed by human annotation errors treated as “ground truth”. Instead, we introduced novel evaluation metrics that are independent of the annotation accuracy. Using grain boundary detection in UO2 as a case study, we found that our approach led to a 64% increase in the total number of grains detected, which acts a robust and holistic measure of model performance on the TEM dataset used in this work. Finally, we showed that model self-confidence is only achieved through transfer learning and fine-tuning of very deep layers.
@@ 1 Introduction
Nanoscale defects, such as grain boundaries, precipitates, and dislocations, play a critical role in controlling the properties and functionality of solid-state materials. Transmission electron microscopy (TEM) has become an irreplaceable tool for investigating these defects, owing to its ultrahigh spatial resolution (sub-Angstrom) [18]. Furthermore, recent advances in faster electron detection and data processing have enabled a big-data approach to characterization techniques such as in-situ TEM and 4D-scanning transmission electron microscopy (4D-STEM) [8]. Terabytes of data can be created in a single-hour session during 4D-STEM or in-situ TEM experiments [5]. However, extracting meaningful insights from the TEM images has quickly become an enormous bottleneck, since the traditional method of manual image analysis is time-consuming, subject to human bias, and cannot scale with the growing data volume [11]. Therefore, developing high-quality automated approaches for TEM image analysis is of paramount importance.
Since the discovery of convolutional neural networks (CNN), machine learning (ML) models have been able to outperform not only traditional computer vision techniques, but even human abilities on certain image analysis tasks [7]. In the context of TEM, binary segmentation utilizing models from the U-Net family enables pixel-level classification of defect structures, which is especially important for identifying continuous defects like grain boundaries and phase interfaces [10]. However, comparing applications to more typical datasets from other sciences, the performance of CNN models for TEM image analysis remains inferior. For instance, most of the literature involving U-Net based models, including the seminal paper [9] and popular derivatives like U-Net++ [23], is focused on medical imaging [12]. With these datasets, it is very common to see F1-scores in the range 0.85-0.95. Similar performance can also be seen in applications like forestry [15], crack detection [6], satellite imaging [2], and plant disease [16]. TEM applications, however, usually report lower F1-scores in range 0.5–0.8 [1][17].
@@
To address these challenges, we explored the use of transfer learning by leveraging advanced ML models pre-trained on ImageNet. In particular, we investigated promising combinations of a pre-trained EfficientNet encoder with a U-Net++ decoder, which helped achieve the best performance in 2023 and 2024 for edge detection in regular optical image datasets [21][22]. We also introduced two novel metrics, prediction certainty and prediction abundance, describing the ability of a model to make predictions with high class probabilities. Together, they define a model’s self-confidence and are independent of ground truth accuracy. We demonstrated that U-Net performance can be substantially improved using a pre-trained encoder, fine-tuning of deep layers, and L2-regularization to control overfitting. Although F1-scores remained limited due to ground truth flaws and uncertainty, improved model self-confidence led to a 64% increase in the number of detected grains for the dataset tested, even surpassing human abilities and significantly enhancing the reliability of defect statistics.
As a practical example, we applied our workflow to a TEM image dataset of nanocrystalline UO2 samples used to study grain growth as function of temperature and heavy ion irradiation dose [20]. UO2 is the primary fuel used in current nuclear reactors, and many of its key material properties, such as thermal conductivity, fission gas retention, and fracture toughness, are governed by its grain size. Therefore, it is critical to establish reliable correlations between grain size of UO2 and its irradiation conditions for predicting the nuclear fuel performance. The accuracy of said correlations principally depends on the quality of grain statistics, which in turn requires large and representative datasets. As such, hundreds of TEM images like those in Fig. 1 were collected at various dose levels and temperatures. Processing these images by hand is infeasible, so ML models are needed to automate segmentation of grain boundaries. To broaden the scope of this study, our methodology was also applied to additional defect types, as discussed in Sec. B of supplementary materials.
@@ 2.1 Grain Boundary Dataset
The dataset used in this work consists of bright-field (BF) TEM images of nanocrystalline UO2 samples. Details about fabrication, in-situ irradiation, and TEM imaging can be found in [20][19]. The training dataset was created by selecting 14 images from a variety of experimental conditions and resizing them from $\times4096$ to $\times1024$. The grain boundaries were then labeled using Inkscape with a thickness of 5 pixels. Pixels belonging to a grain boundary were given a value of 1, while pixels belonging to the grain interior were given a value of 0. A one-to-one mapping is ultimately constructed between TEM and annotated images. Examples can be seen in Fig. 1.
Small dataset size can lead to significant variation in model performance when different validation and test hold-out sets are used. To improve their representation of the training set, each image was quartered into four $\times512$ images, and 5-fold cross-validation was used. This divides the dataset into five non-overlapping validation sets of equal size, training a model for each one with the same hyperparameters. Statistics can then be calculated for performance metrics, providing a more robust estimate. This also means a test set is not necessary, further improving validation set representation. To improve model performance, training images were augmented eightfold using 90° rotations and flipping over horizontal and vertical axes. Each model thus used a different combination of 360 training images and 11 validation images.
@@ 2.2 Model Architectures and Hyperparameters
To improve U-Net performance, the encoder is replaced with EfficientNetB7, the largest model from the EfficientNet-v1 family [14], and the decoder has been replaced with U-Net++. Most promisingly, EfficientNetB7-U-Net++ has been used as the base network of the state-of-the-art edge detectors UAED [21] and MUGE [22], which are applicable to grain boundary segmentation. Furthermore, EfficientNetB7 has been pre-trained on ImageNet [4] to ease training and boost encoder performance for small datasets. The benefits of transfer and fine-tuning are evaluated in Sec. 4. EfficientNetB7 and U-Net++ were also isolated to show their respective contributions, leading to the four architectures investigated in this study, as summarized in Fig. 3: UNet (b), UNet++ (c), Eff-UNet (d), and Eff-UNet++ (e).
The Adam optimizer was used with a global learning rate of 1e-4, as preliminary testing showed that this provided sufficient convergence speed, while having minimal loss fluctuation. It was also found that batch size minimally affected validation loss due to the already large image size ($\times512$), so a batch size of 4 was used to improve batch representation while still offering frequent weight updates. For the yellow convolution blocks in Fig. 2b-e, each Conv2D layer learned $2^{(}i+6)$ filters, where $i=0,1,2,3,4$ refers to the encoder resolution level. ReLU activation was used for all layers except the output layer which used Sigmoid activation. To suppress overfitting, L2-regularization was used to penalize large weights in favor of a more distributed network. The total loss can be written symbolically as
[…]
where BCE is binary cross-entropy loss, commonly used in binary segmentation, $\mathbf{w}$ is a vector of model weights, and $\lambda$ is the L2-regularization strength. In preliminary testing, it was found $\lambda$ can have substantial effects on model output, specifically on model self-confidence. Given that the architectures in Fig. 2 have different sizes and connections, and thus very different $\mathbf{w}$, a coarse grid search was performed to determine the optimal $\lambda$. This grid search simultaneously provided the basis for model comparison.
@@ 2.3 Performance Metrics
To evaluate model performance, classification metrics, i.e., precision, recall, and F1-score, were used with a binarization threshold of 0.5. Two novel metrics, prediction certainty and abundance, were also used to evaluate model self-confidence, defined as the model’s ability to be certain of its own predictions regardless of classification performance. As shown in Fig. 3, the proposed metrics operate on raw prediction image histograms, where certainty measures the ability of a model to predict high class probabilities, and abundance measures the proportion of said pixels that are positive predictions. A model with high certainty tends to avoid uncertain predictions where a pixel could belong to either a grain boundary or non-grain boundary with nearly equal probability. A model with high abundance tends to predict more grain boundaries overall, which is important for a class-imbalanced dataset. We describe model self-confidence comparatively as the possession of both certainty and abundance, which, importantly, do not rely on the existence of a ground truth.
The final metric, detected defect count, was used to provide an unsupervised, holistic measure of model usefulness, as it operates on post-processed model predictions. For the grain boundary dataset, the post-processor Convex Hull Approximate Contour (CHAC) algorithm developed recently in previous work [19] was used. CHAC uses traditional computer vision techniques to identify approximately closed and convex contours in model predictions as grains, based on thermodynamic principles for grain growth. As such, models that can produce segmentation maps leading to a greater number of grains identified by CHAC are more useful since grain size statistics will be more reliable and accurate. Detected defect count therefore implicitly depends on all ML metrics evaluated, as a model must have good classification performance and self-confidence to predict well-defined and crisp grain boundaries. Finally, because each TEM image, and thus each validation set, has a different number of grains, the total number of defects were tabulated across the entire dataset, made possible by $k$-fold cross-validation.
To examine model performance more comprehensively, additional supervised holistic metrics describing sample accuracy, including intersection-over-union and L1 distance metrics, were employed with the results discussed in Sec. A of supplementary materials.
@@ 3.1 Visual Comparison of Predictions
Initial insights into model performance can be gained by examining raw model predictions visually. A representative validation prediction is shown for all architectures and L2-regularization strengths in Fig. 4. When no regularization was used, all architectures appeared to perform similarly. However, upon increasing $\lambda$, UNet and UNet++ responded quite differently than Eff-UNet and Eff-UNet++. The latter architectures exhibited greater tolerance to regularization, as over-regularization occurred at much higher $\lambda$ values. Furthermore, predictions made near $\lambda$=1e-3 demonstrated significant quality improvements, where models clearly predict higher class probabilities, and thus have greater certainty. This has the primary benefit of reducing grain boundary noise in binarized predictions, as predicted probabilities are farther from the threshold (0.5), and are therefore less susceptible to random noise. Finally, unlike changing the encoder, there was no obvious difference between using U-Net and U-Net++ decoders.
3.2 Performance Metrics Analysis
@@
Fig. 5 summarizes the six metrics used to evaluate model architectures with L2-regularization strengths ($\lambda$). Like the visual results in Fig. 4, architectures behave similarly when $\lambda$=0, other than Eff-UNet and Eff-UNet++ having marginally lower precision and recall. As $\lambda$ is increased, recall, certainty, abundance, and detected defect count reached their maximum at $\lambda$=1e-4 for UNet and UNet++, and between $\lambda$=1e-3 and $\lambda$=5e-4 for Eff-UNet and Eff-UNet++. Coincidentally, precision reached its minimum at these points. Furthermore, the following observations can be made: (1) F1-score showed little-to-no improvement despite considerable improvement to detected defect count, (2) there is a large divergence between U-Net and EfficientNet encoder architectures, with the latter showing much greater certainty, abundance, and detected defect counts in Figs. 5d-f, and (3) given the width of the error bars, there is no practical benefit from using the U-Net++ decoder.
@@
Most significantly, by using EfficientNet and L2-regularization, detected defect count increased by 64% over unregularized UNet, even outperforming humans by about 300 grains. This is clearly captured by improvements to certainty and abundance, but the conventional metrics in Figs. 5a-c do not explicitly show this, other than a 7% improvement to recall. This highlights the value of model self-confidence and holistic metrics like detected defect count. As discussed in the next section, the F1-score may not be entirely reliable for images that have higher ambiguity and the human annotation contains more errors.
@@
Figure 5: Metrics obtained from 5-fold cross-validation for $\lambda$ optimization. Figs. a-c are classification metrics precision, recall, and F1-score respectively. Figs. d-e are proposed metrics certainty, abundance, and detected defect count respectively. Error bars correspond to $\pm 1\sigma$
@@ 3.3 Importance of L2-Regularization
As shown in Fig. 5c, the F1-score is limited to 0.75, regardless of architecture or regularization strength. To investigate this limitation, classification performance was visualized in Fig. 6 using a technique we named Confusion Matrix Color Mapping (CM2). Here, every pixel in a binarized prediction image is colored according to its classification outcome in the binary confusion matrix, i.e., green for true positive, blue for false positive, and red for false negative pixels. Importantly, orange arrows indicate examples of grain boundaries which were well-supported by the TEM image and were detected by the Eff-UNet model, but not present in the human annotation. In other words, the human annotator missed these grain boundaries. Annotation errors act as a significant limitation of classification performance where good model predictions are incorrectly penalized.
To minimize loss, a model attempts to learn a set of operations that replicate the human annotated grain boundaries. However, when true grain boundaries are mistakenly missing in the annotation, the model cannot rationalize these inconsistencies, potentially damaging model self-confidence. Any attempt to exactly fit these annotation errors acts as a contradictory force to its typically well-supported operations. If annotation errors are interpreted by the model as high-level features with complex semantic content, the model could begin to overfit by trying to learn these features. The benefits of L2-regularization then become obvious: by constraining model complexity, it discourages fitting such annotation errors, and can instead help the model prioritize simpler, more reliable cues, thus promoting model self-uncertainty.
@@
The theory above is well supported by quantitative results in Fig. 5. Without regularization, all four architectures have relatively high precision and low recall, meaning they guess less but are usually correct, and have very low prediction certainty and abundance. This indicates that they are skeptical; models prefer to guess less grain boundaries, but those that are guessed tend to be more correct. However, applying optimal L2-regularization, the architectures shift toward low precision and high recall, where they now guess more often and are less concerned with correctness. Certainty and abundance also increase substantially, suggesting less influence from human annotation errors and more confidence with their predictions. Among all four architectures, Eff-UNet and Eff-UNet++ show the most significant improvement to prediction quality with detected defect count reaching the highest observed levels, likely because of the greater capacity and redundancy supplied by EfficientNetB7.